Fincorp will pay a year-end dividend of $3.60 per share, which is expected to grow at a rate of 3% for the indefinite future. The discount rate is 12%.

a. What is the stock selling for? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b. If earnings are $3.80 a share, what is the implied value of the firm’s growth opportunities? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

a) D1= 3.6 G= 3% and Re= 12%

P0= 3.6*1.03/(0.12-0.03)

= 41.2

b) PV of growth oppurtunity= E1/(re-g) - E1/re

3.8/0.09 - 3.8/0.12

42.22- 31.67= 10.5533